OpenStudy (anonymous):

If this item was made of crystal, it would need to be packed very carefully in a box with soft filling or packing material all around. If you use a cube shaped box with a height of 16cm, what is the volume you need to fill with packing material?

OpenStudy (anonymous):

A Cube's forumla is V=a^3 a is the only number you got (16) your equation would be V= (16)^3 V= 4096cm^3 hope this helps!